# MsEdF: A Multi-stream Encoder-decoder Framework for Remote Sensing Image Captioning

### v3

@@ 1 Introduction
Remote sensing image captioning (RSIC) refers to the task of automatically generating descriptive sentences from high‑resolution remote sensing images. These images often contain complex spatial layouts, overlapping semantic objects, and ambiguous visual features that challenge traditional representation methods. A widely used approach is the encoder and decoder framework, where a convolutional neural network (CNN) extracts semantic features and a sequential model, typically an RNN or LSTM, generates textual descriptions [18, 16]. To improve performance, previous work has used fully convolutional encoders [20], summarization-driven training [21], and SVM based decoders [7]. More recent efforts include improving the decoder embedding layer with TextGCN embeddings [4] and dynamic graph refinement using CLIP-based semantic and spatial fusion [14].
Most existing approaches for RSIC still rely on conventional single-stream encoder and decoder architectures, which often struggle to model the complex spatial composition and semantic ambiguity in remote sensing images [7, 8]. These models typically adopt a single CNN encoder, such as ResNet [4], InceptionNet [8], or VGGNet [16], without exploring how encoder diversity impacts captioning performance [3]. However, single encoders are limited in capturing both global context and fine-grained object-level details due to constrained receptive fields and uniform feature hierarchies [4], which reduces the expressiveness of extracted visual features.
@@ 3 Proposed Method
The overall architecture is illustrated in Figure˜1. We adopt an encoder-decoder framework, where the encoder ($F$) consists of two distinct CNNs whose outputs are concatenated and processed through a linear transformation layer ($L_{1}$). For decoding, a tokenized input sequence passes through an embedding layer ($E$) and then into a stack of Gated Recurrent Units (GRU) ($G$) to model temporal dependencies. The outputs from $L_{1}$ and $G$ are concatenated and passed through two additional linear layers ($L_{2}$ and $L_{3}$), with $L_{3}$ serving as a softmax layer to generate word probabilities.
The overall architecture of our work is shown in Figure˜1. We follow an encoder-decoder architecture, where the encoder ($F$) encompasses two separate CNNs, whose outputs are concatenated and processed by a linear transformation layer ($L_{1}$). In the decoder, an input sequence in tokens is stored in an embedding layer ($E$) followed by series of GRU cells ($G$) to capture temporal relationships. The outputs of $L_{1}$ and $G$ are concatenated together, which is then used as input to two more linear layers $L_{2}$, and another layer $L_{3}$ serving as a softmax to generate word probabilities.
@@ 5.3 Visual Interpretation
We present qualitative results from test images in Figure˜3. Here we have compared the proposed model (MsEdF) with ground-truth captions (Gold) and conventional encoder-decoder models (ConvNext and ResNet). These examples highlight the cases where the proposed approach produces captions that are semantically rich. Several examples (Figures˜3a, 3b and 3c) show cases where the baseline model produces incorrect or misleading captions, whereas the proposed model generates accurate and contextually relevant descriptions. In Figure˜3a, ConvNext and ResNet misclassify the image as industrial area. In Figure˜3b, ConvNext misclassifies the image as resort and ResNet as school. In Figure˜3c, ConvNext and ResNet misclassify the image as mobile home park. Other example (Figures˜3d, 3e and 3f) highlights that, although both models produce generally correct captions, the proposed approach incorporates extra visual or semantic details that enhance the overall quality of the caption. In Figure˜3d, while all models identify mountain, the proposed model further recognizes snow. In Figure˜3e, while all models identify park, the proposed model further recognizes lake and road, enhancing the semantic richness of the caption. In Figure˜3e, while all models identify ocean or river  [^1] Visually, it is challenging to distinguish whether it is an ocean or an river; therefore, both interpretations are accepted. , the proposed model also recognizes runway with marking lines.
@@ 5.4 Subjective Evaluation
Table 13: Subjective Evaluation of Different CNNs on Three Datasets (in %)

### v4

@@ 1 Introduction
Remote sensing image captioning (RSIC) refers to the task of automatically generating descriptive sentences from high‑resolution remote sensing images. A widely adopted approach for generating accurate captions from remote sensing images is the encoder–decoder framework. Here a convolutional neural network (CNN) extracts semantic features and a sequential model, typically an RNN or LSTM, generates textual descriptions [18, 16]. Most existing approaches for RSIC still rely on conventional single-stream encoder and decoder architectures, which often struggle to model the complex spatial composition and semantic ambiguity in remote sensing images [7, 8]. These models typically adopt a single CNN encoder, such as ResNet [4], InceptionNet [8], or VGGNet [16], without exploring how encoder diversity impacts captioning performance [3]. However, single encoders are limited in capturing both global context and fine-grained object-level details due to constrained receptive fields and uniform feature hierarchies [4], which reduces the expressiveness of extracted visual features.
@@ 3 Proposed Method
We follow an encoder-decoder architecture, where the encoder ($F$) encompasses two separate CNNs, whose outputs are concatenated and processed by a linear transformation layer ($L_{1}$). In the decoder, an input sequence in tokens is stored in an embedding layer ($E$) followed by series of GRU cells ($G$) to capture temporal relationships. The outputs of $L_{1}$ and $G$ are concatenated together, which is then used as input to two more linear layers $L_{2}$, and another layer $L_{3}$ serving as a softmax to generate word probabilities. The overall architecture of our work is shown in Figure˜1.
@@ 5.3 Visual Interpretation
We present qualitative results from test images in Figure˜3. Here we have compared the proposed model (MsEdF) with ground-truth captions (Gold) and conventional encoder-decoder models (ConvNext and ResNet). These examples highlight the cases where the proposed approach produces captions that are semantically rich. Several examples (Figures˜3a, 3b and 3c) show cases where the baseline model produces incorrect or misleading captions, whereas the proposed model generates accurate and contextually relevant descriptions. In Figure˜3a, ConvNext and ResNet misclassify the image as industrial area. In Figure˜3b, ConvNext misclassifies the image as resort and ResNet as school. In Figure˜3c, ConvNext and ResNet misclassify the image as mobile home park. Other example (Figures˜3d, 3e and 3f) highlights that, although both models produce generally correct captions, the proposed approach incorporates extra visual or semantic details that enhance the overall quality of the caption. In Figure˜3d, while all models identify mountain, the proposed model further recognizes snow. In Figure˜3e, while all models identify park, the proposed model further recognizes lake and road, enhancing the semantic richness of the caption. In Figure˜3f, while all models identify ocean or river (Visually, it is challenging to distinguish whether it is an ocean or an river; therefore, both interpretations are accepted), the proposed model also recognizes runway with marking lines.
@@ 5.4 Subjective Evaluation
Table 13: Subjective Evaluation of Different Encoder-decoder models on Three Datasets (in %)
@@
5.5 Error Analysis
During visual interpretation of different models, several common issues were identified in the generated captions. The most common problem is misclassification [16], where similar objects are confused. For example: industrial buildings as resort or school (Figure˜3b), dense residential area as mobile home parks (Figure˜3c), river as pond and farmland as bareland. A more subtle error involves contextual confusion, where the presence of an object affects the overall interpretation of the image. For example, buildings within the airport image causes a misclassification of the image as industrial area (Figure˜3a). In some images, the main object is missing. An example is the absence of a runway (Figure˜3f). Minor object omissions also occur, such as missing houses in meadow, green trees in bridge image, snow in mountains (Figure˜3d)), and lake in park (Figure˜3e). Additional issues include attribute misclassification (such as count, color), falsely adding non-existent objects, and incomplete sentence generation.